# Generalized Simple Graphical Rules for Assessing Selection Bias

### v3

@@ Abstract
Selection bias is a major obstacle toward valid causal inference in epidemiology. Over the past decade, several simple graphical rules based on causal diagrams have been proposed as the sufficient identification conditions for addressing selection bias and recovering causal effects. However, these simple graphical rules are usually coupled with specific identification strategies and estimators. In this article, we show two important cases of selection bias that cannot be addressed by these simple rules and their estimators: one case where selection is a descendant of a collider of the treatment and the outcome, and the other case where selection is affected by the mediator. To address selection bias in these two cases, we construct identification formulas by the g-computation and the inverse probability weighting (IPW) methods based on single-world intervention graphs (SWIGs). They are generalized to recover the average treatment effect by adjusting for post-treatment upstream causes of selection. We propose two IPW estimators and their variance estimators to recover the average treatment effect in the presence of selection bias in these two cases. We conduct simulation studies to verify the performance of the estimators when the traditional crude selected-sample analysis returns erroneous contradictory conclusions to the truth.
@@ 1 Introduction
Selection bias is one of the fundamental obstacles to causal inference. The concept of selection bias encompasses various forms stemming from different mechanisms for individuals being selected into the study or analysis [1]. The complexity of these selection mechanisms poses difficulty in a structural definition of selection bias, which has been continually evolving over the past two decades [2, 3, 4, 5]. Recently, a refined definition of selection bias that unifies different types of selection bias was proposed [3] by considering the true causal effect in the population prior to the selection process as the reference. This definition can be decomposed into two types: type 1 selection bias due to restricting to at least one level of a collider or its descendant, and type 2 selection bias due to restricting to at least one level of an effect measure modifier.[3]
@@
With a clearer understanding of selection bias in place, recent methodological developments have also been dedicated to recovering causal effects in the presence of selection bias. Specifically, based on the fundamentals of causal diagrams, including directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs), simple graphical rules were proposed to account for selection bias, alone or possibly combined with other issues (e.g., confounding bias or missing data) [6, 7, 8, 9].
However, there are still important cases of selection bias that cannot be addressed by these existing graphical rules. For example, in the graphical rules proposed by Mathur and Shpitser, one premise is that post-treatment variables cannot be conditioned on in their selected-sample analysis estimator [8]. In addition, they focus on the conditional average treatment effect (CATE) rather than the average treatment effect (ATE, which is arguably more widely used). Some other existing graphical rules do not prohibit adjustments for post-treatment variables, but they are rather stringent to avoid confounding bias [10, 7, 6].
@@
In some cases, the covariate information available to address selection bias may only include post-treatment variables [11]. In these cases, the identification of ATE may still be possible despite the rejection by the existing graphical rules. This highlights the importance of applying fundamental graphical rules to search for identification strategies. For example, approaches such as inverse probability weighting (IPW) methods and g-computation are powerful tools for recovering treatment effects if external information beyond the selected sample is available [1, 7, 11].
@@
In this article, we discuss two important cases of selection bias unaddressed by existing graphical rules. The first one represents the selection bias where a descendant of a collider of the exposure and the outcome is being conditioned on. The other is an example where sample selection is affected by the mediator. We prove that the average treatment effect (ATE) in both cases can be recovered by the g-computation or IPW methods if the external information on the covariates is available in the general population. The conditions for proving these identification strategies are provided and can also be used for addressing selection bias in similar cases. We propose two specific IPW estimators and their variance estimators for the ATE. We conduct simulation studies to verify the statistical properties of the proposed IPW estimators. In our hypothetical simulation settings, we show that the conventional crude complete-case analysis using the selected sample yields conclusions contradictory to the truth and can mislead decision-making.
@@ 2 Two cases of selection bias unaddressed by simple graphical rules
To detect and control for selection bias, researchers usually consider some simple graphical rules along with a specific class of identification formulas, and then define ‘identifiability’ (or ‘recoverability’) in a narrow sense of whether the treatment effect of interest can be identified by that class of formulas [10, 7, 6, 8]. Simple graphical rules are usually sufficient conditions for verifying that the treatment effect can be identified (or recovered) by a particular class of identification formulas. If these conditions are violated, it does not rule out the possibility that the effect of interest can still be identified by other formulas. In this section, we show two important cases of selection bias that cannot be addressed by the previously proposed graphical rules. The identification of treatment effects for similar causal structures by modified formulas will be discussed in later sections.
@@ 2.1 Setting
Consider $n$ individuals (hereafter called “unit”) in a randomized clinical trial whose outcomes are observed only among selected units. The selection is directly affected by a post-treatment variable observable among all units. Specifically, let $A_{i}$ be the binary treatment for unit $i$, $Y_{i}$ be the binary outcome, and $S_{i}$ be the selection. There are no other variables as confounders of $A_{i}$ and $Y_{i}$. We define the direct upstream cause of the selection to be $L_{i}$, which can be either a common effect of the treatment and the outcome (Case 1 in Fig. 1), or a mediator of the treatment on the outcome (Case 2 in Fig. 1). When a unit is selected, i.e. $S_{i}=1$, we observe all variables $(A_{i},Y_{i},L_{i})$. When a unit is not selected, i.e. $S_{i}=0$, we observe its direct upstream cause $L_{i}$ along with the knowledge of $A_{i}$ or the rule of treatment assignment $p(A_{i}=1)$ while the outcome $Y_{i}$ remains unobserved. We denote potential responses to a particular treatment $A_{i}=a_{i}$ by $(Y_{i}(a_{i}),L_{i}(a_{i}),S_{i}(a_{i}))$, which are affected by the treatment in both cases. Under the consistency assumption, $(Y_{i}(a_{i}),L_{i}(a_{i}),S_{i}(a_{i}))$ = $(Y_{i},L_{i},S_{i})$.
Both Case 1 and Case 2 are important as they represent scenarios that may occur in real-world settings. For example, consider Case 1, where the study is a randomized trial of a prophylactic medication for malaria among susceptibles. The treatment $A$ is a binary variable indicating assignment to the medication or not. The outcome of interest $Y$ is the incidence of malaria or not. The incidence of fever or not, $L$, directly affects the receipt of the blood test for malaria or not, $S$, for the outcome to be detected. It was suggested that the selection bias in this exact causal graph cannot be addressed to recover the ATE by the selection-backdoor criteria [6].
In Case 2, the study is a randomized trial of a vaccine for an infectious disease with multiple endpoints. The treatment $A$ indicates assignment to the vaccine or not. The outcome $Y$ is the symptom of having a headache or not. The outcome is mediated by the incidence of infection, $L$, which also affects whether a unit reports the outcome status, $S$.
@@
$A$$Y$$L$$S$$A$ $a$ $Y(a)$$L(a)$$S(a)$$A$$L$$Y$$S$$A$ $a$$L(a)$$Y(a)$$S(a)$Case 1Case 2DAGSWIG
Figure 1: Causal directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs) for Case 1 and Case 2 of sample selection directly affected by a post-treatment variable $L$. When a unit is selected into the sample (i.e. $S_{i}=1$), we observe all variables $(A_{i},Y_{i},L_{i})$. When a unit is not selected (i.e. $S_{i}=0$), external information on $L_{i}$ among unselected units, along with the knowledge of $A_{i}$ or the rule of treatment assignment $p(A_{i}=1)$, might still be available, but the outcome $Y_{i}$ is unobserved. In Case 1, the direct upstream cause of selection, $L$, is the common effect (i.e., collider) of both the treatment $A$ and the outcome $Y$. In Case 2, the direct upstream cause of selection, $L$, is the mediator of the treatment $A$ on the outcome $Y$.
2.2 The failure by existing simple graphical rules
We focus on three existing simple graphical rules (i.e., the selection-backdoor criterion [6], the generalized adjustment criterion [7], and Mathur and Shpitser’s graphical rules [8]) proposed over the past decade. In Appendix A, we show that the two cases as discussed above violate these existing simple graphical rules, and the identification by their corresponding formulas is not guaranteed.
3 Identification by the fundamentals of graphical rules
In this section, we first introduce a g-computation formula that can be used to recover the ATE by conditioning on post-treatment variables [11]. Then we propose the inverse probability weighting (IPW) method to recover the ATE and address selection bias in the two cases discussed. The identification results by the two proposed methods are generalized under specific graphical rules to allow (i) observed confounders for the relationships between the treatment and post-treatment variables, (ii) unobserved confounders for the relationships between post-treatment variables and the outcome, and (iii) selection that is directly affected by the treatment.
@@ 3.1 G-computation formulas
In Breskin et al.[11], a special-case g-formula conditioning on a post-treatment variable $Z$ was derived for identifying the ATE in a practical example of randomized controlled trials, demonstrating the utility of SWIGs. We extend their results by allowing the existence of confounders in an observational setting.
@@
Given a SWIG $\mathcal{G}(a)$ containing the random treatment component $A$, the fixed treatment component $a$, the post-treatment variables $Z(a)$, the selection variable $S(a)$, the outcome variable $Y(a)$, the observed pre-treatment variables $X$, and the unobserved pre-treatment variables $U$, the ATE is identified by
$\begin{split}\delta_{\text{ATE},g}=\sum_{x}(\sum_{z}&\mathbb{E}[Y|A=1,Z=z,S=1,X=x]p(Z=z|A=1,X=x)-\\ &\mathbb{E}[Y|A=0,Z=z,S=1,X=x]p(Z=z|A=0,X=x))p(X=x),\end{split}$ (1)
@@
$\begin{split}&(C1)\quad Y(a)\perp\!\!\!\!\perp S(a)|Z(a),X\\ &(C2)\quad Y(a)\perp\!\!\!\!\perp A|Z(a),S(a),X\\ &(C3)\quad Z(a)\perp\!\!\!\!\perp A|X.\end{split}$ (2)
Intuitively, the first condition assumes that all paths between $Y(a)$ and $S(a)$ are blocked by $Z(a)$ and $X$. This motivates the adjustment of some post-treatment variables $Z(a)$ if it helps separate the outcome from the selection. The next two conditions assume that the effect of the treatment on both $Y(a)$ and $Z(a)$ are unconfounded by conditioning on different sets of variables.
Therefore, despite being motivated by an ad-hoc estimator for a special case, the g-formula (1) will recover the ATE when some post-treatment variables can be conditioned on to address selection bias. The conditions can be viewed as derived graphical rules in this specific context. The example by Breskin et al. [11] is a special case of our theorem (where they denote the post-treatment variable by $W$ instead of $L$) by the following corollary.
Corollary 1 (Identification of the ATE by the g-formula given SWIG [11]).
@@
$\begin{split}\delta_{\text{ATE},g}=\sum_{w}&\mathbb{E}[Y|A=1,W=w,S=1]p(W=w|A=1)-\\ &\mathbb{E}[Y|A=0,W=w,S=1]p(W=w|A=0),\end{split}$
i.e. $\delta_{\text{ATE}}=\delta_{\text{ATE},g}$, for the SWIG in Breskin et al.[11].
Taking $Z=L$, our Case 2 can be considered as a modification of the graph in Breskin et al. [11] where the derivation of the formula still applies. Moreover, it is straightforward to verify that the conditions (2) are satisfied in both of our cases by the SWIGs in Fig. 1. We have the following corollary as examples.
@@
For estimation and inference by the g-formulas, plug-in estimators can be applied and the variances can be derived using the delta method or bootstrap [14].
@@
Figure 2: Causal directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs) for Case 1 and Case 2 of sample selection directly affected by a post-treatment variable $L$ with observed covariates $X=(X_{1},X_{2},X_{3})$ and unobserved covariates $U$. When a unit is selected (i.e. $S_{i}=1$), we observe all variables $(X_{i},A_{i},Y_{i},L_{i})$. When a unit is not selected (i.e. $S_{i}=0$), external information on $(X_{i},A_{i},L_{i})$ among unselected units might still be available. Case 1: The direct upstream cause of selection, $L$, is the common effect (i.e., collider) of the treatment $A$ and the outcome $Y$. Case 2: The direct upstream cause of selection, $L$, is the mediator on the causal pathway from treatment $A$ to the outcome $Y$.
@@ 3.2 Inverse probability weighting
Inverse probability weighting (IPW) methods intuitively create a pseudo-population where the distributions of covariates are balanced across treatment groups to fix the endogeneity of treatment and recover its effects [15, 16]. They have also been recognized and applied for the generalizability and transportability of study results to a target population of interest that is different from the study sample [17, 18, 19, 20, 21, 22].
@@
Given a SWIG $\mathcal{G}(a)$ containing the random treatment component $A$, the fixed treatment component $a$, the post-treatment variables $Z(a)$, the selection variable $S(a)$, the outcome variable $Y(a)$, the observed pre-treatment variables $X$, and the unobserved pre-treatment variables $U$, the ATE is identified by
$\begin{split}\delta_{\text{ATE},\text{IPW}}=\mathbb{E}[\frac{ASY}{p(S=1|Z,X,A)p(A=1|X)}]-\mathbb{E}[\frac{(1-A)SY}{p(S=1|Z,X,A)p(A=0|X)}],\end{split}$ (3)
@@
$\begin{split}&(C1)\quad Y(a)\perp\!\!\!\!\perp S(a)|Z(a),X\\ &(C2)\quad(Y(a),S(a))\perp\!\!\!\!\perp A|Z(a),X\\ &(C3)\quad Z(a)\perp\!\!\!\!\perp A|X\\ \end{split}$ (4)
The first condition assumes that all paths between $Y(a)$ and $S(a)$ are blocked by $Z(a)$ and $X$. The next two conditions assume that the treatment and the potential responses of post-treatment variables (including the outcome, covariates, and selection) can be sequentially separated. The second condition for the IPW method (4) is slightly different from that for the g-formula (2). The requirement by the IPW method that $S(a)$ should be conditionally separated from $A$ is due to the inclusion of the inverse probability weighting model for the selection to recover the ATE.
The ATE can be recovered through the IPW method in all causal graphs in Breskin et al.[11], Fig. 1, and Fig. 2. They are special cases of Theorem 2 by the following corollaries.
Corollary 4 (Identification of the ATE by the IPW method given SWIG [11]).
@@
i.e. $\delta_{\text{ATE}}=\delta_{\text{ATE},\text{IPW}}$, for the SWIG in Breskin et al.[11].
@@
The treatment-selection confounders $X_{3}$ cannot be omitted by the IPW method for recovering the ATE because the model for the selection deems it necessary to collect and adjust for these variables for correct causal model specification. This is different from the requirements by the g-formula for the same graph.

### v4

# On the Graphical Rules for Recovering the Average Treatment Effect Under Selection Bias

@@ Abstract
Selection bias is a major obstacle toward valid causal inference in epidemiology. Over the past decade, several graphical rules based on causal diagrams have been proposed as the sufficient identification conditions for addressing selection bias and recovering causal effects. However, these simple graphical rules are typically coupled with specific identification strategies and estimators. In this article, we show two important cases of selection bias that cannot be addressed by these existing simple rules and their estimators: one case where selection is a descendant of a collider of the treatment and the outcome, and the other case where selection is affected by the mediator. To address selection bias and recover average treatment effect in these two cases, we propose an alternative set of graphical rules and construct identification formulas by the g-computation and the inverse probability weighting (IPW) methods based on single-world intervention graphs (SWIGs). We conduct simulation studies to verify the performance of the estimators when the traditional crude selected-sample analysis (i.e., complete-case analysis) yields erroneous conclusions contradictory to the truth.
@@ 1 Introduction
Selection bias is one of the fundamental obstacles to causal inference. The concept of selection bias encompasses various forms stemming from different mechanisms for individuals being selected into the study or analysis [hernan2004structural]. The complexity of these selection mechanisms poses difficulty in a structural definition of selection bias, which has been continually evolving over the past two decades [hernan2017invited, lu2022toward, kenah2023potential, lu2024evolution]. Recently, a refined definition of selection bias that unifies different types of selection bias was proposed [lu2022toward] by considering the true causal effect in the population prior to the selection process as the reference. This definition can be decomposed into two types: type 1 selection bias due to restricting to at least one level of a collider or its descendant, and type 2 selection bias due to restricting to at least one level of an effect measure modifier.[lu2022toward]
@@
With a clearer understanding of selection bias in place, recent methodological developments have been dedicated to recovering causal effects in the presence of selection bias. Specifically, aided by causal diagrams, including directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs)[richardson2013single, bezuidenhout2025swigs], several graphical rules were proposed to assess and account for selection bias, either alone or in combination with other challenges (e.g., confounding bias or missing data) [bareinboim2014recovering, correa2018generalized, mathur2023simple, schnell2024graphical].
While these graphical rules provide valuable insights, they are typically paired with specific identification strategies or estimators. For example, under the graphical rules proposed by Mathur and Shpitser, one key premise is that post-treatment variables cannot be conditioned on when conducting selected-sample analysis (i.e., complete-case analysis) [mathur2023simple]. In addition, their focus is on the conditional average treatment effect (CATE), rather than the average treatment effect (ATE) which is arguably more widely used. By contrast, epidemiologic analyses have frequently relied on inverse probability of selection or censoring weights to address different forms of selection bias (e.g., due to non-random sample selection or differential loss to follow-up). [hernan2004structural, howe2016selection] These approaches leverage covariate information (potentially including post-treatment variables) from unselected or censored individuals, rather than relying solely on selected-sample analysis. Consequently, the graphical rules proposed by Mathur and Shpitser may not be universally applicable. Other existing graphical rules, while not prohibiting adjustment for post-treatment variables, impose stringent restrictions intended to avoid confounding bias [correa2017causal, correa2018generalized, bareinboim2014recovering].
@@
Nonetheless, some important cases of selection bias fall outside the scope of existing graphical rules. In certain settings, for example, the covariate information available to address selection bias may consist solely of post-treatment variables [breskin2018practical]. In such cases, identification of causal effects (particularly the ATE) may still be possible using estimators such as g-computation [breskin2018practical, zivich2025gcomp], even though identification would be ruled out by existing graphical rules. This underscores the value of applying fundamental graphical principles to guide the search for appropriate identification strategies. For instance, inverse probability weighting (IPW) and g-computation remain powerful tools for recovering treatment effects when external information beyond the selected sample is available [hernan2004structural, correa2018generalized, breskin2018practical].
@@
In this article, we examine two important cases of selection bias that are not addressed by existing graphical rules. The first arises when conditioning on a descendant of a collider between the exposure and the outcome. The second occurs when sample selection is affected by a mediator. We show that in both cases, the ATE can be recovered using g-computation or IPW, provided that external covariate information from the unselected sample is available. We formally state the conditions under which these identification strategies hold, and demonstrate how they can be applied to address selection bias in related settings. In addition, we propose two IPW estimators for the ATE, along with corresponding variance estimators, and evaluate their statistical properties through simulation studies. In our hypothetical simulation settings, we find that conventional complete-case analyses based solely on the selected sample yield conclusions that contradict the true causal effects and risk misleading decision-making.
@@ 2 Two cases of selection bias unaddressed by simple graphical rules
To detect and control selection bias, researchers often rely on simple graphical rules in conjunction with a specific class of identification formulas, and define identifiability (or recoverability) narrowly as whether the treatment effect of interest can be identified within that class [correa2017causal, correa2018generalized, bareinboim2014recovering, mathur2023simple]. These graphical rules typically provide sufficient conditions for establishing that the treatment effect is identifiable (or recoverable) using a given set of formulas. However, failure to satisfy these conditions does not imply that the effect is unidentifiable; rather, it indicates that identification may require alternative strategies or estimators. In this section, we present two important cases of selection bias that fall outside the scope of previously proposed graphical rules. Identification of treatment effects for the related causal structures using modified formulas is discussed in later sections.
@@ 2.1 Setting
Consider $n$ individuals (hereafter called “unit”) in a randomized clinical trial whose outcomes are observed only among selected units. The selection is directly affected by a post-treatment variable observable among all units. Specifically, let $A_{i}$ be the binary treatment for unit $i$, $Y_{i}$ be the binary outcome, and $S_{i}$ be the selection. There are no other variables as confounders of $A_{i}$ and $Y_{i}$. We define the direct upstream cause of the selection to be $L_{i}$, which can be either a common effect of the treatment and the outcome (Case 1 in Fig. 1), or a mediator of the treatment on the outcome (Case 2 in Fig. 1). When a unit is selected, i.e. $S_{i}=1$, we observe all variables $(A_{i},Y_{i},L_{i})$. When a unit is not selected, i.e. $S_{i}=0$, we possibly observe a part or none of the variables. Under the consistency assumption, $(Y_{i}(a_{i}),L_{i}(a_{i}),S_{i}(a_{i}))$ = $(Y_{i},L_{i},S_{i})$.
@@
AAYYLLSSAAaaY(a)Y(a)L(a)L(a)S(a)S(a)AALLYYSSAAaaL(a)L(a)Y(a)Y(a)S(a)S(a)Case 1Case 2DAGSWIG
Figure 1: Causal directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs) for Case 1 and Case 2 of sample selection directly affected by a post-treatment variable $L$. When a unit is selected into the sample (i.e. $S_{i}=1$), we observe all variables $(A_{i},Y_{i},L_{i})$. When a unit is not selected (i.e. $S_{i}=0$), we may observe a part or none of the variables. In Case 1, the direct upstream cause of selection, $L$, is the common effect (i.e., collider) of both the treatment $A$ and the outcome $Y$. In Case 2, the direct upstream cause of selection, $L$, is the mediator of the treatment $A$ on the outcome $Y$.
2.2 Limitations of existing simple graphical rules
We focus on three existing graphical rules proposed over the past decade: the selection-backdoor criterion [bareinboim2014recovering], the generalized adjustment criterion [correa2018generalized], and the graphical rules of Mathur and Shpitser [mathur2023simple]. In Appendix A, we demonstrate that the two cases discussed above violate these rules, and that identification by their corresponding formulas is therefore not guaranteed.
3 Identification by the fundamentals of graphical rules with external data
In this section, we first introduce a g-computation formula that can recover the ATE by conditioning on post-treatment variables [breskin2018practical] using external data (i.e., covariate information from the unselected sample). We then propose IPW estimators to recover the ATE and address selection bias in the two cases described above. The identification results from these methods are further generalized under specific graphical rules to allow (i) observed confounders between the treatment and post-treatment variables, (ii) unobserved confounders between post-treatment variables and the outcome, and (iii) selection processes directly affected by the treatment.
@@ 3.1 G-computation formulas
In Breskin et al. [breskin2018practical], a special-case g-formula conditioning on a post-treatment variable $L$ was derived to identify the ATE in a randomized controlled trial example, illustrating the utility of SWIGs. We extend their results to observational settings by allowing for the presence of confounders. Our identification results hold when external information on covariates among unselected units is available, together with knowledge of $A_{i}$ or the treatment assignment mechanism $p(A_{i}=1)$, while permitting the outcome $Y_{i}$ to remain unobserved.
@@
Given a SWIG $\mathcal{G}(a)$ containing the random treatment component $A$, the fixed treatment component $a$, the post-treatment variables $L(a)$, the selection variable $S(a)$, the outcome variable $Y(a)$, the observed pre-treatment variables $X$, and the unobserved pre-treatment variables $U$, the ATE is identified by
$\begin{split}\delta_{\text{ATE},g}=\sum_{x}(\sum_{l}&\mathbb{E}[Y|A=1,L=l,S=1,X=x]p(L=l|A=1,X=x)-\\ &\mathbb{E}[Y|A=0,L=l,S=1,X=x]p(L=l|A=0,X=x))p(X=x),\end{split}$ (1)
@@
$\begin{split}&(C1)\quad Y(a)\perp\!\!\!\!\perp S(a)|L(a),X\\ &(C2)\quad Y(a)\perp\!\!\!\!\perp A|L(a),S(a),X\\ &(C3)\quad L(a)\perp\!\!\!\!\perp A|X.\end{split}$ (2)
Intuitively, the first condition assumes that all paths between $Y(a)$ and $S(a)$ are blocked by $L(a)$ and $X$. This motivates the adjustment of some post-treatment variables $L(a)$ if it helps separate the outcome from the selection. The next two conditions assume that the effect of the treatment on both $Y(a)$ and $L(a)$ are unconfounded by conditioning on different sets of variables.
Therefore, despite being motivated by an ad-hoc estimator for a special case, the g-formula (1) will recover the ATE when some post-treatment variables can be conditioned on to address selection bias. The conditions can be viewed as derived graphical rules in this specific context. The example by Breskin et al. [breskin2018practical] is a special case of our theorem by the following corollary.
Corollary 1 (Identification of the ATE by the g-formula given SWIG [breskin2018practical]).
@@
i.e. $\delta_{\text{ATE}}=\delta_{\text{ATE},g}$, for the SWIG in Breskin et al.[breskin2018practical].
Our Case 2 can be considered as a modification of the graph in Breskin et al. [breskin2018practical] where the derivation of the formula still applies. Moreover, it is straightforward to verify that the conditions (2) are satisfied in both of our cases by the SWIGs in Fig. 1. We have the following corollary as examples.
@@
For estimation and inference by the g-formulas, plug-in estimators can be applied and the variances can be derived using the delta method or bootstrap [imai2010general].
@@
Figure 2: Causal directed acyclic graphs (DAGs) and single-world intervention graphs (SWIGs) for Case 1 and Case 2 of sample selection directly affected by a post-treatment variable $L$ with observed covariates $X=(X_{1},X_{2},X_{3})$ and unobserved covariates $U$. When a unit is selected (i.e. $S_{i}=1$), we observe all variables $(X_{i},A_{i},Y_{i},L_{i})$. When a unit is not selected (i.e. $S_{i}=0$), we may observe a part or none of the variables. Case 1: The direct upstream cause of selection, $L$, is the common effect (i.e., collider) of the treatment $A$ and the outcome $Y$. Case 2: The direct upstream cause of selection, $L$, is the mediator on the causal pathway from treatment $A$ to the outcome $Y$.
@@ 3.2 Inverse probability weighting
Inverse probability weighting (IPW) methods intuitively create a pseudo-population where the distributions of covariates are balanced across treatment groups to fix the endogeneity of treatment and recover its effects [rosenbaum1983central, austin2015moving]. They have also been recognized and applied for the generalizability and transportability of study results to a target population of interest that is different from the study sample [cole2010generalizing, stuart2011use, lesko2017generalizing, dahabreh2019generalizing, dahabreh2020extending, degtiar2023review].
@@
Given a SWIG $\mathcal{G}(a)$ containing the random treatment component $A$, the fixed treatment component $a$, the post-treatment variables $L(a)$, the selection variable $S(a)$, the outcome variable $Y(a)$, the observed pre-treatment variables $X$, and the unobserved pre-treatment variables $U$, the ATE is identified by
$\begin{split}\delta_{\text{ATE},\text{IPW}}=\mathbb{E}[\frac{ASY}{p(S=1|L,X,A)p(A=1|X)}]-\mathbb{E}[\frac{(1-A)SY}{p(S=1|L,X,A)p(A=0|X)}],\end{split}$ (3)
@@
[…]
The first condition assumes that all paths between $Y(a)$ and $S(a)$ are blocked by $L(a)$ and $X$. The next two conditions assume that the treatment and the potential responses of post-treatment variables (including the outcome, covariates, and selection) can be sequentially separated. The second condition for the IPW method (4) is slightly different from that for the g-formula (2). The requirement by the IPW method that $S(a)$ should be conditionally separated from $A$ is due to the inclusion of the inverse probability weighting model for the selection to recover the ATE.
The ATE can be recovered through the IPW method in all causal graphs in Breskin et al.[breskin2018practical], Fig. 1, and Fig. 2. They are special cases of Theorem 2 by the following corollaries.
Corollary 4 (Identification of the ATE by the IPW method given SWIG [breskin2018practical]).
@@
i.e. $\delta_{\text{ATE}}=\delta_{\text{ATE},\text{IPW}}$, for the SWIG in Breskin et al.[breskin2018practical].
@@
The treatment-selection confounders $X_{3}$ cannot be omitted by the IPW method for recovering the ATE because the selection model requires adjustment for all variables that influence sample selection, including $X_{3}$, to ensure correct causal model specification. This requirement differs from that of the g-formula for the same graph.